# Geodesic Calculus on Latent Spaces

### v1

@@ Abstract.
Figure 6. Left: Visualization of sample points in latent space (projected from $\mathbb{R}^{10}$ into $\mathbb{R}^{3}$ based on a PCA) and, starting from a fixed point in a fixed direction, linear extrapolation (red) and geodesic extrapolation with $\mathcal{W}_{{\mathcal{M}}}$ (green). Right: Corresponding decoded sequences.
@@
Next, we consider image data of a rotating three-dimensional object as proposed as an example by Braunsmann et al. (2024). We use their regularized autoencoder, minimizing a loss function that promotes the embedding to be as isometric and as flat as possible. The data ${\mathcal{X}}\subset{\mathcal{M}}\subset\mathbb{R}^{128\times 128\times 3}$ consists of RGB images showing a toy cow model (Crane et al., 2013) from varying viewpoints. Training the regularized autoencoder requires computing distances and averages between dataset samples. Each image $x$ corresponds to a specific rotation $r_{x}\in\mathrm{SO}(3)$ which allows to define these distances and averages. We use the publicly available pretrained autoencoder with $l=16$-dimensional latent space, for details on other parameters and settings see section˜A.6. A PCA projection on three dimensions of the resulting latent manifold is visualized in fig.˜7 (left), where the PCA analysis shows that the encoder uses six dimensions for the embedding.
@@
As the encoder is regularized to be near-isometric, we use $\mathcal{W}_{\text{E}}(z,\tilde{z})=|z-\tilde{z}|^{2}$ in our geodesic calculus. In fig.˜7, linear and geodesic interpolation as well as extrapolation in latent space are shown.
@@
Finally, our discrete geodesic interpolation can be performed even if the latent manifold $\mathcal{Z}$} is merely represented by a distance function $d\colon\mathbb{R}^{l}\to\mathbb{R}$ with ${\mathcal{Z}}=\{{z}\in\mathbb{R}^{l}\mid d(z)=0\}$. However, $d$ is non-differentiable on $\mathcal{Z}$}. Hence, instead of the augmented Lagrange algorithm (9)-(10) we use the classical penalty method to minimize the discrete path energy ${\mathcal{E}}^{K}(\mathbf{{z}})$ subject to the constraints $d({z}_{k})=0$. Computing the exponential map is not possible with missing normal information, though.
For example, Pose-NDF (Tiwari et al., 2022) provides a neural distance function to a manifold of plausible human poses. They use a quaternion representation of the SMPL body model by Loper et al. (2015), resulting in $l=84$ dimensions. Comparisons with linear interpolation in these coordinates show that our approach computes geodesics on the manifold of plausible poses, whereas linear interpolation produces self-intersecting poses that lie off the manifold, see fig.˜8. An additional example is provided in section˜A.7.
@@ Appendix A Appendix
    $\mathcal{W}_{\mathrm{E}}({z}_{0},{z}_{1})\coloneqq|{z}_{1}-{z}_{0}|^{2}$
@@
In practice, we use a slightly more advanced version of the Augmented Lagrangian method following the algorithm described by Nocedal & Wright (2006, Chapter 17), which we adapt to our setting in algorithm˜1. Compared to the simplified version in the main text, we do not update the Lagrange multiplier and the penalty parameter in every iteration, but only depending on how well the constraint is already fulfilled. For the inner optimization problem, we use the BFGS method from SciPy (Virtanen et al., 2020). As the initial path we choose the path $\mathbf{{z}}_{0}=\bigl(z_{0}=z_{0},\ldots,z_{\left\lfloor K/2\right\rfloor}=z_{0},\;z_{\left\lfloor K/2\right\rfloor+1}=z_{K},\ldots,z_{K}=z_{K}\bigr)$ that remains constant and jumps directly from the given starting point $z_{0}$ to the endpoint $z_{K}$ at the middle time point. We set the initial Lagrange multiplier $\Lambda_{i0}=0$ for $i=1,\ldots,K$ and $\alpha=2$. The final tolerance $\eta^{\ast}$ for the constraint is problem-dependent and depends on the minimum values of $\zeta_{\sigma}$. In practice, a good rule-of-thumb is to choose $\eta^{\ast}$ as $K$ times the mean value on the embedded data samples $\eta^{\ast}\approx\frac{K}{|{\mathcal{X}}|}|\zeta_{\sigma}(\phi({\mathcal{X}}))|$.
@@
We train a fully connected neural network with ELU activations (Clevert et al., 2015) by minimizing the loss functional ˜12 using the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-3}$ and a weight decay of $10^{-5}$. The layer dimensions depend on the examples. The batch size to evaluate the integral over $\mathcal{Z}$} is $128$ in all our examples. To approximate the inner integral, we sample a single point $y_{z}$ from $f_{\sigma}$ for each data sample $z$ and optimization step.
@@
To analyze the denoising loss and the resulting approximate projection $\Pi_{\sigma}$ for different parameters, we use the toy torus model to allow comparison with a ground truth projection and keep the evaluation visually tractable. In practice, only approximations of the latent manifold $\mathcal{Z}$} are available. To study the denoising property, we generate a noisy torus surface and train a projection with different values for $\sigma$, treating the noisy surface as $\mathcal{Z}$}. We then visualize the image $\Pi_{\sigma}({\mathcal{Z}})$ under the projections. As expected, a larger value of $\sigma$ leads to a stronger smoothing effect, see fig.˜9.
For a point cloud without noise, a small parameter $\sigma$ leads to a higher accuracy of $\Pi_{\sigma}$ close to the surface but larger errors at certain distances, as those points are rarely seen in training, see fig.˜10 (left). We further evaluate in fig.˜10 the stability of the optimization, showing that the approximation error decreases for increasing point cloud size, increasing network architectures, and decreasing noise levels.
@@
All computations in this stage follow established numerical algorithms: For the Riemannian center of mass, the logarithms, and the exponential map, we use the methods introduced by Heeren et al. (2014) with time resolution $K=8$. For the tangent PCA, we use the representation using edge lengths and dihedral angles as described by Sassen et al. (2020). We used the first two components for the example in fig.˜1 and the first ten components for the example in figs.˜4 and 11. To this end, we employed their publicly available C++ implementation (Heeren & Sassen, 2020).
@@
The training objective combines the reconstruction loss with the isometry loss introduced by Braunsmann et al. (2021) (see below). To generate training samples $\mathcal{X}$}, we proceed as follows: First, we uniformly sample points within a hyperball of the linear tangent subspace. The size of the hyperball was chosen based on the norms of the projections of the input Riemannian logarithms onto the subspace. Second, for each such point, we sample a nearby point using a normal distribution with small variance centered on the first point. Finally, we apply the discrete exponential map to the two points to obtain a pair of points on the submanifold. We discard any pairs where at least one shape exhibits self-intersections to preserve physical plausibility. For the isometry loss, we compute the distance between the two shapes in a pair by computing discrete geodesics and taking their length. For the example in fig.˜1, we drew $10000$ pairs this way, and, for the example in figs.˜4 and 11, we drew $100000$ pairs. All these computations were performed with the same setup as described above.
@@
In fig.˜11, we provide additional comparisons between linear interpolation in the latent space and geodesic interpolation using our learned implicit representation.
Figure 11. Two comparisons between linear interpolation (red) and geodesic interpolation using a learned projection on the embedded submanifold of discrete shells (green).
@@
We train the projection on the embedded samples by embedding the full set of training data and sampling one point per resulting vMF distribution. We use a fully connected network with layers (10, 64, 64, 64, 10) and choose $\sigma=0.05$.
In fig.˜12, we provide an additional example of geodesic interpolation. Moreover, as further variant, in this figure we also show a path computed using the pullback metric $\mathcal{W}_{{\mathcal{M}}}$ but without using the implicit representation $\zeta_{\sigma}$ as constraint.
Figure 12. Same as fig.˜5 for a different pair of endpoints. Left: Visualization of sample points in latent space (projected from $\mathbb{R}^{10}$ into $\mathbb{R}^{3}$) and computed paths between two points via linear interpolation (red), geodesic interpolation with $\mathcal{W}_{\text{E}}$ (yellow), unconstrained interpolation with $\mathcal{W}_{{\mathcal{M}}}$ (blue), and geodesic interpolation with $\mathcal{W}_{{\mathcal{M}}}$ (green). Right: Corresponding decoded sequences.
@@
Each image $x$ corresponds to a specific rotation $r_{x}\in\mathrm{SO}(3)$. Hence, a distance between the images can be defined as $\mathrm{dist}_{\mathcal{M}}(x,y)=\arccos(r_{x}\cdot r_{y})$ and geodesic averages ${\mathrm{av}}_{\mathcal{M}}(x,y)$ as renderings of the object with the mean rotation between $r_{x}$ and $r_{y}$. The autoencoder is trained with tuples of nearby points $(x,y)\in{\mathcal{X}}_{\epsilon}$, where ${\mathcal{X}}_{\epsilon}\subset\{(x,y)\in{\mathcal{M}}\times{\mathcal{M}}\mid\,\mathrm{dist}_{\mathcal{M}}(x,y)\leq\epsilon\}$. The regularization loss ${\mathcal{J}}_{\mathrm{reg}}$ for the encoder is given by
@@
Figure˜13 shows an additional example using the neural distance function Pose-NDF (Tiwari et al., 2022) as manifold representation, see section˜5.4. We compare linear interpolation in the quaternion representation of the SMPL body model used by Tiwari et al. (2022) and geodesic interpolation on the manifold corresponding to the approximate zero-level set.

### v2

@@ Abstract.
Figure 7. Decoded sequences starting from a fixed point in a fixed direction; linear extrapolation (red) and geodesic extrapolation with $\mathcal{W}_{{\mathcal{M}}}$ (green).
@@
Next, we consider image data of a rotating three-dimensional object as proposed as an example by Braunsmann et al. (2024). We use their regularized autoencoder, minimizing a loss function that promotes the embedding to be as isometric and as flat as possible. The data ${\mathcal{X}}\subset{\mathcal{M}}\subset\mathbb{R}^{128\times 128\times 3}$ consists of RGB images showing a toy cow model (Crane et al., 2013) from varying viewpoints. Training the regularized autoencoder requires computing distances and averages between dataset samples. Each image $x$ corresponds to a specific rotation $r_{x}\in\mathrm{SO}(3)$ which allows to define these distances and averages. We use the publicly available pretrained autoencoder with $l=16$-dimensional latent space, for details on other parameters and settings see section˜B.3. A PCA projection on three dimensions of the resulting latent manifold is visualized in fig.˜8 (left), the PCA analysis shows that the encoder uses six dimensions.
@@
As the encoder is regularized to be near-isometric, we use $\mathcal{W}_{\text{E}}(z,\tilde{z})=|z-\tilde{z}|^{2}$ in our geodesic calculus. In fig.˜8, linear and geodesic interpolation as well as extrapolation in latent space are shown.
@@
Finally, our discrete geodesic interpolation can be performed even if the latent manifold $\mathcal{Z}$} is merely represented by a distance function $d\colon\mathbb{R}^{l}\to\mathbb{R}$ with ${\mathcal{Z}}=\{{z}\in\mathbb{R}^{l}\mid d(z)=0\}$. However, $d$ is non-differentiable on $\mathcal{Z}$}. Hence, instead of the augmented Lagrange algorithm (10)-(11) we use the classical quadratic penalty method to minimize the discrete path energy ${\mathcal{E}}^{K}(\mathbf{{z}})$ subject to the constraints $d({z}_{k})=0$. Computing the exponential map is not possible with missing normal information, though.
For example, Pose-NDF (Tiwari et al., 2022) provides a neural distance function to a manifold of plausible human poses. They use a quaternion representation of the SMPL body model by Loper et al. (2015), resulting in $l=84$ dimensions. Comparisons with linear interpolation in these coordinates show that our approach computes geodesics on the manifold of plausible poses, whereas linear interpolation produces self-intersecting poses that lie off the manifold, see fig.˜9. An additional example is provided in section˜B.4.
@@ Appendix A Method details
    $\mathcal{W}_{\mathrm{E}}({z},\tilde{z})=|\tilde{z}-{z}|^{2}$
@@
In practice, we use a slightly more advanced version of the Augmented Lagrangian method following the algorithm described by Nocedal & Wright (2006, Chapter 17), which we adapt to our setting in algorithm˜1. Compared to the simplified version in the main text, we do not update the Lagrange multiplier and the penalty parameter in every iteration, but only depending on how well the constraint is already fulfilled. For the implementation of the minimization in ˜10 we implemented two versions, one using the LBFGS method from the PyTorch optimization package and a NumPy version using the BFGS method from SciPy (Virtanen et al., 2020). As the initial path we choose the path $\mathbf{{z}}_{0}=\bigl(z_{0}=z_{0},\ldots,z_{\left\lfloor K/2\right\rfloor}=z_{0},\;z_{\left\lfloor K/2\right\rfloor+1}=z_{K},\ldots,z_{K}=z_{K}\bigr)$ that remains constant and jumps directly from the given starting point $z_{0}$ to the endpoint $z_{K}$ at the middle time point. This is preferable to linear interpolation for initialization, as it ensures that the initial paths lies on or close to the manifold. We set the initial Lagrange multiplier $\Lambda_{i0}=0$ for $i=1,\ldots,K$ and $\alpha=2$. The final tolerance $\eta^{\ast}$ for the constraint is problem-dependent and depends on the minimum values of $\zeta_{\sigma}$. In practice, a good rule-of-thumb is to choose $\eta^{\ast}$ as $K$ times the mean value on the embedded data samples $\eta^{\ast}\approx\frac{K}{|{\mathcal{X}}|}|\zeta_{\sigma}(\phi({\mathcal{X}}))|$. The initial penalty parameter should be chosen sufficiently large to ensure that the first approximate solution remains close to the manifold.
@@
We train a fully connected neural network with ELU activations (Clevert et al., 2015) by minimizing the loss functional ˜13 using the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-3}$ and a weight decay of $10^{-5}$. The layer dimensions depend on the examples. The batch size to evaluate the integral over $\mathcal{Z}$} is $128$ in all our examples. To approximate the inner integral, we generate a single sample $y_{z}=z+\epsilon$ where $\epsilon\sim f_{\sigma}$ for each data sample $z$ and optimization step.
@@
To analyze the denoising loss and the resulting approximate projection $\Pi_{\sigma}$ for different parameters, we use the toy torus model to allow comparison with a ground truth projection and keep the evaluation visually tractable. In practice, only approximations of the latent manifold $\mathcal{Z}$} are available. To study the denoising property, we generate a noisy torus surface and train a projection with different values for $\sigma$, treating the noisy surface as $\mathcal{Z}$}. We then visualize the image $\Pi_{\sigma}({\mathcal{Z}})$ under the projections. As expected, a larger value of $\sigma$ leads to a stronger smoothing effect, see fig.˜10.
For a point cloud without noise, a small parameter $\sigma$ leads to a higher accuracy of $\Pi_{\sigma}$ close to the surface but larger errors at certain distances, as those points are rarely seen in training, see fig.˜11 (left). We further evaluate in fig.˜11 the stability of the optimization, showing that the approximation error decreases for increasing point cloud size, increasing network architectures, and decreasing noise levels.
@@ Appendix B Experiment details
All computations in this stage follow established numerical algorithms: For the Riemannian center of mass, the logarithms, and the exponential map, we use the methods introduced by Heeren et al. (2014) with time resolution $K=8$. For the tangent PCA, we use the representation using edge lengths and dihedral angles as described by Sassen et al. (2020). We used the first two components for the example in fig.˜1 and the first ten components for the example in figs.˜5 and 12. To this end, we employed their publicly available C++ implementation (Heeren & Sassen, 2020).
@@
The training objective combines the reconstruction loss with the isometry loss introduced by Braunsmann et al. (2021) (see below). To generate training samples $\mathcal{X}$}, we proceed as follows: First, we uniformly sample points within a hyperball of the linear tangent subspace. The size of the hyperball was chosen based on the norms of the projections of the input Riemannian logarithms onto the subspace. Second, for each such point, we sample a nearby point using a normal distribution with small variance centered on the first point. Finally, we apply the discrete exponential map to the two points to obtain a pair of points on the submanifold. We discard any pairs where at least one shape exhibits self-intersections to preserve physical plausibility. For the isometry loss, we compute the distance between the two shapes in a pair by computing discrete geodesics and taking their length. For the example in fig.˜1, we drew $10000$ pairs this way, and, for the example in figs.˜5 and 12, we drew $100000$ pairs. All these computations were performed with the same setup as described above.
@@
In fig.˜12, we provide additional comparisons between linear interpolation in the latent space and geodesic interpolation using our learned implicit representation.
Figure 12. Two comparisons between linear interpolation (red) and geodesic interpolation using a learned projection on the embedded submanifold of discrete shells (green). Away from self intersections linear and geodesic interpolation are qualitatively similar indicating that the manifold is flat in these regions (bottom).
@@
We train the projection on the embedded samples by embedding the full set of training data and sampling one point per resulting vMF distribution. We use a fully connected network with 4 layers of 64 intermediate dimensions and Leaky ReLU activation functions, and we choose $\sigma=0.05$.
In fig.˜13, we provide an additional example of geodesic interpolation. Moreover, as further variant, in this figure we also show a path computed using the pullback metric $\mathcal{W}_{{\mathcal{M}}}$ but without using the implicit representation $\zeta_{\sigma}$ as constraint.
Figure 13. Same as fig.˜6 for a different pair of endpoints. Left: Visualization of sample points in latent space (projected from $\mathbb{R}^{10}$ into $\mathbb{R}^{3}$) and computed paths between two points via linear interpolation (red), geodesic interpolation with $\mathcal{W}_{\text{E}}$ (yellow), unconstrained interpolation with $\mathcal{W}_{{\mathcal{M}}}$ (blue), and geodesic interpolation with $\mathcal{W}_{{\mathcal{M}}}$ (green). Right: Corresponding decoded sequences.
@@
Each image $x$ corresponds to a specific rotation vector $r_{x}$. Hence, a distance between the images can be defined as $\mathrm{dist}_{\mathcal{M}}(x,y)=\arccos(r_{x}\cdot r_{y})$ and geodesic averages ${\mathrm{av}}_{\mathcal{M}}(x,y)$ as renderings of the object with the mean rotation between $r_{x}$ and $r_{y}$. The autoencoder is trained with tuples of nearby points $(x,y)\in{\mathcal{X}}_{\epsilon}$, where ${\mathcal{X}}_{\epsilon}\subset\{(x,y)\in{\mathcal{M}}\times{\mathcal{M}}\mid\,\mathrm{dist}_{\mathcal{M}}(x,y)\leq\epsilon\}$. The regularization loss ${\mathcal{J}}_{\mathrm{reg}}$ for the encoder is given by
@@
Figure˜14 shows an additional example using the neural distance function Pose-NDF (Tiwari et al., 2022) as manifold representation, see section˜5.4. We compare linear interpolation in the quaternion representation of the SMPL body model used in Tiwari et al. (2022) and geodesic interpolation on the manifold corresponding to the approximate zero-level set.
@@
Appendix C Comparisons
In this section, we compare different numerical approaches to compute geodesics in settings related to the one considered in this article.
[…]
A parametric viewpoint represents a manifold as the image of a coordinate map from a lower-dimensional parameter domain. This viewpoint is taken in many numerical approaches such as the well-established Stochman library Detlefsen et al. (2021). In contrast, we take the viewpoint of an implicit manifold defined as the zero level set of a function in an ambient (latent) space. In both cases, geodesics between fixed endpoints are computed as solutions of a variational problem: minimizing the path energy. In our setting, this minimization is subject to the constraint to stay on the zero level set defined in ˜1 and is incorporated in the time-discrete scheme as described in section˜3. The exponential map (i.e. the geodesic for a prescribed initial point and velocity) corresponds to the solution of an initial value problem for a second-order ordinary differential equation. Instead of employing off-the-shelf PyTorch-based ODE solvers as in the Stochman library, we use a variational ODE solver, which ensures that the discrete path computed by the discrete exponential map well approximates a discrete geodesic between its start and end point. As a toy model, we compute geodesics on the torus ${\mathcal{M}}=\{z=({z}^{1},{z}^{2},{z}^{3})\in\mathbb{R}^{3}\mid 0=\zeta(z):=(\sqrt{({z}^{1})^{2}+({z}^{2})^{2}}-R)^{2}+({z}^{3})^{2}-r^{2}\}$. In our approach, we minimize the discrete path energy
(15) ${\mathcal{E}}^{K}({z}_{0},\ldots,{z}_{K})=K\sum_{k=1,\ldots,K}\|{z}_{k}-{z}_{k-1}\|^{2}$
for prescribed endpoints and under the constraint $\zeta({z}_{k})=0$ for all $k=0,\ldots,K$. For the parametric approach, we parametrize the torus by $0,2\pi)^{2}\ni(\theta_{1},\theta_{2})\mapsto((R+r\cos\theta_{2})\cos\theta_{1},(R+r\cos\theta_{2})\sin\theta_{1},r\sin\theta_{2})\in\mathbb{R}^{3}$, which yields the metric tensor
$G(\theta_{1},\theta_{2})=\begin{pmatrix}\bigl(R+r\cos\theta_{2}\bigr)^{2}&0\\ 0&r^{2}\end{pmatrix}$.
Using the Stochman library Detlefsen et al. (2021), the curve energy induced by this metric is minimized using a cubic spline approximation evaluated at discrete timesteps on the two-dimensional parameter domain. Using our approach with $50$ nodes ($K=49$), the minimization takes 1.89 seconds and yields an approximated geodesic with path energy $4.7889$. To reach comparable accuracy with Stochman, we use a spline with 8 nodes and 50 metric evaluation points and apply 2000 iterations for the optimization, resulting in a path energy of $4.7892$. The minimization takes 2.48 seconds and no qualitative difference between the different discrete geodesics is visible; see fig.˜15 (left).
Using our variational discretization for the exponential map (see section˜3) and considering the first two points of the discrete variational geodesic to define the initial direction, we obtain a discrete exponential map with 50 timesteps in less than a second. It reproduces the discrete variational geodesic; see fig.˜15 (right). Using the tool for the exponential map in the Stochman library, the computation also takes less than one second. However, the result deviates visibly from that of our variational scheme, which is a manifestation of the robustness of our scheme and points to the moderate instability observed when computing exponential maps in negatively curved spaces. In applications to real data sets, we consider robustness to be particularly desirable. Finally, note that our approach of discretizing geodesics is not restricted to implicit manifolds; in fact, it was originally formulated for a parametric viewpoint Rumpf & Wirth (2015).
Figure 15. Left: geodesics computed with our approach (green) and the Stochman library (yellow), right: exponential maps for the initial direction taken from our discrete path energy minimizing geodesic shown for our approach (green) and for the Stochman library tool (yellow).
[…]
In the setting of latent manifolds, one does not have direct access to a continuous metric tensor or a parameterized surface representation, but only to a discrete manifold representation given by samples. Several approaches have been proposed to compute geodesics in this setting. Some methods treat the manifold as a point cloud and rely on graph-based shortest-path algorithms Sorrenson et al. (2025); Mckenzie & Damelin (2019). However, such approaches are generally not suited for computing an exponential map. Arvanitidis et al. (2022) precompute curve energies on a uniform grid in latent space, use graph-based shortest-path methods to obtain an initial geodesic, and subsequently refine it via cubic-spline optimization. This grid-based construction does not scale well with increasing latent dimension, and special treatment is required to connect the start and end points to the computed geodesic. In Arvanitidis et al. (2019), the authors discuss ill-conditioning and failure modes of standard solvers for computing geodesics on learned manifolds. As an alternative, they propose a fixed-point iteration scheme that avoids explicit Jacobian evaluations. A specific discretization of the exponential map is not discussed. In Shao et al. (2018), the exponential map is approximated via parallel transport. We will revisit several of these methods in the comparisons presented in section˜C.3 and section˜C.4.
[…]
In Arvanitidis et al. (2022), the Kullback–Leibler (KL) divergence or equivalently its second-order approximation (14), is used as an approximation of the Fisher information metric on the latent manifold of variational autoencoders (VAEs). To compute geodesics on the latent manifold, this metric is not sufficient. Indeed, information on the actual geometry of the latent manifold as a submanifold of the latent space is needed. As a remedy, Arvanitidis et al. (2022) extend the metric to be very costly away from the latent manifold, such that shortest paths avoid crossing this outside region. In principle, this extension can be done in different ways; Arvanitidis et al. (2022) choose to decode points off the latent manifold as probability distributions with a much larger variance, dubbed “uncertainty regularization”, and then use the above-described KL divergence-based Riemannian metric on the full latent space. The authors calibrate the decoder in a post-processing step via an estimation of distances of latent points to the latent data manifold. They approximate these distances via Euclidean K-means clustering and choose the distance to the nearest cluster center Arvanitidis et al. (2022)[Appendix C.2]. Based on these distances, they apply a sigmoid function to interpolate between the decoder variance and a large fallback variance. Then, geodesics can be computed either by first discretizing the manifold on a grid in latent space, as described in Arvanitidis et al. (2022), or by directly minimizing the curve energy. Note that, just like in our method, this approach requires establishing a representation of the latent manifold, i.e., a distance function, whereas we use a projection operator onto the latent manifold. The subsequent translation of the distance function into a Riemannian metric on the full latent space is solely for the purpose of computing geodesics via unconstrained minimization. Then, geodesics are actually allowed to leave the latent manifold and will typically only stay in its vicinity. Thus, it is conceptually cleaner to directly compute geodesics constrained to the latent manifold, which is the approach we pursue. Furthermore, a distance function is in general less robust to compute than a projection: while the former is nonsmooth for latent manifold generalization with locally varying codimension, the latter is not. Finally, finding the nearest cluster center for each point in latent space to approximate the distance function does not scale well with latent manifold dimension, which is why, in our approach, we instead leverage the power of deep learning to learn a projection.
To compute geodesics, a pullback metric based on the mean and variance of the decoder is used in Arvanitidis et al. (2018), together with an extension to the full latent space using a regularized decoder variance that is large away from the data support. For training, a radial-basis-function network (RBF) Broomhead & Lowe (1988) is employed. Furthermore, a Euclidean $k$-nearest-neighbor graph with edge weights given by Riemannian distances is constructed. Then, at query time, the endpoints are projected onto the graph, a discrete shortest path is extracted, and the resulting polyline is smoothed and interpolated with a time-parameterized cubic spline to obtain an initial continuous geodesic approximation.
A comparison of those two approaches with our method is given in fig.˜16 for a toy distribution. To this end, we train a VAE that decodes into normal distributions with different means and variances and learn the projection onto the latent manifold, which takes approximately 12 minutes. Based on this, we compute geodesics with our method using the Euclidean metric in latent space (4 seconds), the pullback of the Euclidean metric of decoded samples (30 seconds), and for the KL-divergence-based metric (1 minute). For the algorithm accompanying Arvanitidis et al. (2022), the calibration with sigmoid functions takes 5 seconds, and the minimization of the path energy takes 10 seconds. However, the resulting geodesic leaves the data manifold, and the optimization does not converge. In the framework of Arvanitidis et al. (2018) with the radial-basis-function network, the accompanying code takes 21 seconds for calibration and 45 seconds to compute geodesics using the pullback of the mean and variance. While for low-dimensional latent manifolds the calibration of the decoder is faster than learning our projection, these approaches are limited to VAEs and to models without fixed decoder variance. In general, the optimization is harder, together with the previously mentioned conceptual disadvantages.
Figure 16. Top left: 3D input point cloud; bottom left: our learned representation of the 2D learned latent manifold and a visualization of the calibrated covariances using the respective method; right: latent manifold with geodesics computed using different methods.
[…]
There is a line of work that aims at defining density-based metrics on a data distribution that are cheap in areas of high data density and expensive otherwise. This allows to compute trajectories that follow the data distributions. Examples include Fermat distances Trillos et al. (2024), estimating local covariance matrices by fitting normal distributions Arvanitidis et al. (2016), and score-based approaches in the context of diffusion models Yu et al. (2025); Lobashev et al. (2025b).
We compare our method with an approach that fits locally adaptive normal distributions (LAND model) Arvanitidis et al. (2016), implemented using the Stochman library Detlefsen et al. (2021). Rather than fitting full normal distributions, the method restricts itself to computing local samples with diagonal covariance matrices for the local data distribution around each point, and then uses the associated Mahalanobis inner product as a Riemannian metric, weighted by the inverse data point density. We consider two settings: first, the original application to point-cloud-represented data distributions, and second, a latent manifold example. Let us remark that although our method is designed for latent manifolds, it can be applied in both scenarios.
First, we show a comparison on a toy data distribution in fig.˜17, where the data consist of a noisy sampling of a half-circle. Shortest path computations lead to qualitatively similar results for our approach and for the LAND model. We observe that our method performs better in the case of a larger distance between the chosen endpoints and leads to a more stable exponential computation.
Figure 17. Top row from left to right: level sets of our learned representation function, five geodesics with randomly sampled start and end points, a geodesic between endpoints of the half-circle, and the exponential map with initial velocity tangential to the central half-circle. Bottom row: comparison with LAND on the same set of experiments.
In principle, the LAND model can be used in latent spaces with custom metrics different from the Euclidean one. However, as reported in Arvanitidis et al. (2022), this is computationally demanding, as it requires repeated computations of Riemannian exponential maps and logarithms. To cope with this drawback, the authors propose to precompute these maps on a grid in latent space, which is naturally restricted to low-dimensional latent spaces. For a comparison with our method, we therefore consider the example of the close-to-isometric embedding from section˜5.3 with a low-dimensional latent space in order to be able to directly apply the method with the Euclidean metric; see fig.˜18. The computed geodesics of the LAND model are of acceptable quality but show some artifacts, indicating that the discrete geodesic paths leave the neighborhood of the latent manifold. In fact, we emphasize that the model is designed for volumetric data of varying density rather than for manifolds with non-vanishing codimension.
For the experiment in fig.˜17, we use $\sigma=0.04$ for learning the projection, a network with four fully connected layers, 64 intermediate dimensions, and ELU activation functions. The dataset contains 6000 points. Training up to the selected epoch takes approximately ten minutes. Computing all five geodesics with $K=30$ requires less than ten seconds, while the exponential map takes approximately six seconds. For all computations, the initial penalty parameter is set to $\mu_{0}=100$ and the maximum penalty parameter to $\mu_{\text{max}}=1000$. For the LAND model, we use a locality radius of $\sigma=0.04$ and a density regularization of $\rho=10^{-4}$. Computation of all five geodesics with the LAND model takes approximately two minutes, while the exponential map requires about one second. Our parameters for the learned image manifold are as reported in section˜B.3. The computation of a geodesic takes under two minutes for our model and three minutes for the LAND model.
Figure 18. Comparison on the latent manifold from section˜5.3. Top rows (green): geodesics computed with our approach from fig.˜8. Bottom rows (blue): geodesics computed with the LAND model.